OpenStudy (anonymous):

a car moves with constant acceleration along a straight highway. At time t1 = 10s, it is observed to be moving at a speed v1 = 25m/s; at time t2 = 60s, its speed has increased to v2 = 50m/s. Find (1) the acceleration of the car, (ii) the distance traveled in the time interval t1 = 10s to t2 = 60s.

OpenStudy (anonymous):

from what we know, which of the 3 kiematic equations can we use to solve for acceleration, a? we will want to pick an equation that a in it, and that we know all the other variables in it. v=v+at v^2=v^2+2as s = s + vt + (1/2)at^2 can you see which equation to use?